Sunday, 25 January 2009

Maintenance: 00:00 to 06:00 25th Jan 08

Following the air-con problem yesterday it turned out that our elderly server ITO1, which has been performing front line service since we started the company, was not well. We were aware it was pretty heat-sensitive and had already purchased and two installed new fast heat-tolerant replacements with a switch-over scheduled in the next few days. We decided to bring this forward and completed it last night (UK time). ITO1 will now go into museum service and two new servers have taken up the strain.

It is likely, however that we missed an OSM update last night for the second day running last night and are therefore still running with Thursday's OSM dataset.

No comments: